OpenStudy (anonymous):

Find the area of the regular polygon. Give the answer to the nearest tenth. hexagon with a radius of 5 in.

OpenStudy (p0sitr0n):

A = b*h/2 *6 A = b*h * 3 b= r = 5 h = ? a^2 + b^2 = c^2 5^2 - 2.5^2 = h^2 h =4.33 A = 64.95